OpenStudy (anonymous):

so for my maths assignment we are given points to plot then it says this: model 1- Linear -Plot the data given (done) - Use technology to find an approximate linear model of the form A(d) = bd+c, where: b, c are constants A(d)= accuracy as a function of distance, and d= distance from hoop post My thing is about basketballers by the way. Basically, I don't know how to find the values of A, d, b and c because my linear model is y=mx+c not A(d)= bd+c Thankyou!!!

OpenStudy (yrelhan4):

A(d)= bd+c and y=mx+c are the same things. y substituted with a(d), m substituted with b and x substituted with d..

OpenStudy (anonymous):

oh really? Sorry it was a dumb question but thakyou!!

OpenStudy (yrelhan4):

yeah i can even write z=fd+c.. but conventionally we write y=mx+x You're welcome.